Peninsula Candy Company makes three types of candy bars: Chewy, Chunky, and Choco-Lite (Lite). Sales volume for the annual budget is determined by estimating the total market volume for candy bars and then applying the company's prior year market share, adjusted for planned changes due to company programs for the coming year. Volume is apportioned among the three bars based on the prior year's product mix, again adjusted for planned changes for the coming year.


The following are the company budget and the results of operations for July.

 

BudgetChewy
Chunky
Choco-Lite
Total
Sales-units (in thousands)2,000 bars
2,000 bars
4,000 bars
8,000 bars
Sales-dollars (in thousands)
$240



$360



$720



$1,320

Variable costs

180




300




560




1,040

Contribution margin
$60



$60



$160



$280

Manufacturing fixed cost

40




38




60




138

Product margin
$20



$22



$100



$142

Marketing and administrative costs (all fixed)



















55

Operating profit


















$87

























ActualChewyChunkyChoco-LiteTotal
Sales-units (in thousands)1,700 bars
2,000 bars
4,100 bars
7,800 bars
Sales-dollars (in thousands)
$102



$360



$720



$1,182

Variable costs

58




304




564




926

Contribution margin
$44



$56



$156



$256

Manufacturing fixed cost

38




39




59




136

Product margin
$6



$17



$97



$120

Marketing and administrative costs (all fixed)



















44

Operating profit


















$76


 

Industry volume was estimated at 96.5 million bars for budgeting purposes. Actual industry volume for July was 89.745 million bars.


Required:

a. Determine the sales price and sales activity variances.

b. Break down the sales activity variance into the parts caused by industry volume and market share.
Answer #1

step 1) 1,700* (180/2,000)+ 2,000* (300/2,000)+ 4,100* (560/4,000)= 1,027


step 2) budget chart   60/2,000= 0.03     160/4,000= 0.04

Step 3) 1,182- 1,027= 155

(1,700*0.03)+ (2,000*0.03)+ (4,100*0.04)= 275

1,320- 1,040= 280


part A answer) sales price variance   155- 275= 120 U

sales activity variance  275-280= 5 U


part B Step 1) 7,800* (280/8,000)= 273

280* (89,745/ 96,500)= 260.4


market share variance 273- 260.4= 12.6 F

275(from part A)- 260.4= 14.6 F

12.6 or 14.6

Industry variance: 260.4- 280= 19.6 U